Page 1
FOR KARNATAKA CLASS 11 PREPARATION
Karnataka 1st PUC
Model Question Paper
2027 for Mathematics
(Paper 2)
EXAM YEAR TYPE SUBJECT
Karnataka Class 11 2027 Sample Paper Maths
Notes · Sample Papers · Previous Year Papers · Mock Tests
Page 2
m
m c. o
c. o
e m
em l a s
s ag
GOVERNMENTOF KARNATAKA
DEPARTMENT OF SCHOOL EDUCATION (PRE UNIVERSITY)
I P.U.C ANNUAL EXAMINATION 2027
SAMPLE MODEL QUESTION PAPER II
Subject Code: 35
m
c. o
MATHEMATICS
om No. of questions: 47]
c. [TOTAL e m
m s
[Time: 3 Hours] [Max. Marks: 80]
a s e g la
Instructions:
g l a
a
1. The question paper has five Parts namely A, B, C, D and E. Answer all the
parts.
2. PART-A has 15 Multiple choice questions and 5 Fill in the blank questions.
3. For PART-A questions, only the first written answers will be considered for
evaluation.
PART - A
o m
I. Answer ALL the multiple choicequestions
c
. statement is correct?
(15 1 = 15)
e m
s
1. Let A = {1, 2, {3, 4}, 5}. Which of the following
B){3, 4} ∈A la
A){3, 4} ⊂A
2. If (x - 1, y + 2) = (1, 5), then (x,
g
a y) is
C) 3∈A D) 4∈A.
A) (1, 2) B) (2, 3) C) (3, 2) D) (2, 1)
3. The range of f (x) = sin x is
A)[0, ) B) (-, ) C) [-1, 1] D) (-1, 1)].
m
c. o
4. The square roots of – 1
o m m
c
A) 1,-1 B) −𝑖 C) 𝑖 D) 𝑖, −𝑖
. s e
s em A) (-2, )
5. The solution of 3x + 8 > 2, when x is a real number is
g l a
l a a
ag
B) (-, -2) C) (-, -2] D) (-, -1)
6. Match List I with List II
List I List II
a) 5𝐶0 i) 20
b) 5𝑃2 ii) 10
m .
c) 5𝐶2 iii) 1
.co s e m
e m l a
l a s a g
ag For more Question Papers, Sample Papers, Notes & Syllabus visit Page 1 of 4
Page 3
Choose the correct answer from the options given below:
A) a-i, b-ii, c-iii B) a-iii, b-ii, c-i
C) a-ii, b-i, c-iii D) a-iii, b-i, c-ii
7. The number of 2digit even numbers can be formed from the digits 1,
2, 3, 4, 5 if the digits can be repeated
A) 10 B)8 C)120 D)
8. The value of C1 + C2 + C3 + ..... C10 is
10 10 10 10
A)1024 B)1023 (C)1022 (D) 512
9. The 9th term of the sequence defined by 𝒂𝒏 = (− 𝟏)𝒏−𝟏 𝒏𝟑 is
A) 36 B) 39 C) -39 D)-36
10. The angle between the x-axis and the line joining the points (3,–1) and
(4,–2)
A) 450 B) 1350 C) -450 D) -1350
11. Equation of the circle with centre at (-3,2) and radius 4.
A) 𝑥 2 + 𝑦 2 + 6𝑥 − 4𝑦 + 3 = 0 B) 𝑥 2 + 𝑦 2 − 6𝑥 + 4𝑦 − 3 = 0
C) 𝑥 2 + 𝑦 2 − 6𝑥 + 4𝑦 + 3 = 0 D) 𝑥 2 + 𝑦 2 + 6𝑥 − 4𝑦 − 3 = 0
12. The point (-2, -3, -4) lies in the
A) First Octant B)Seventh octant C) Second octant D) Eighth octant
13. lim (x3 – x2 + 1) is
x 1
A) -1 B) 0 C) 1 D) 2
14. The median of 3, 10, 6, 7, 11, 13, 15 is
A) 7 B) 9 C) 10 D) 11
15. Which of the following is incorrect
A) If A and B are any two events then𝑃(𝐴 ∪ 𝐵) = 𝑃 (𝐴) + 𝑃(𝐵) − 𝑃(𝐴 ∩ 𝐵)
B) P(A′ ) = 1 − P(A)
C) P(A ∩ B ′ ) = P(A) − P(A ∩ B)
D)𝑃(𝐴′ ∩ 𝐵′ ) = 1 + 𝑃(𝐴 ∪ 𝐵)
II. Fill in the blanks by choosing the appropriate answer from those given in
the bracket. (0,1, 2, 3, 4, 5) (5 1 = 5)
16. The length of the intervals (-1, 1] is______
17. The solution of −8 ≤ 5𝑥 − 3 < 7 𝑖𝑠a ≤ 𝒙 < 𝒃, 𝒕h𝒆𝒏 𝒂 + 𝒃 is_____
18. The x-intercept of the line 2 x 5 y 10 0 is_____
19. The length of the lotus rectum of the parabola y 2 4 x is_____
20. The derivative of f(x) = 3𝑥 at x =2
PART-B
III. Answer Any Six of the following questions (6 2 = 12)
21. If A= {3, 5, 7, 9, 11}, B= {7, 9, 11, 13}, C= {15, 17} find A∩(B∪C).
For more Question Papers, Sample Papers, Notes & Syllabus visit Page 2 of 4
Page 4
m
m c. o
c. o
e m
s
m 22. In a circle of diameter 40 cm, the length of a chord isla20 cm. Find the length of
se ag
minor arc of the chord
23. Prove that (sin3x+ sin x) sin x + (cos 3x – cos x) cos x = 0
25
24. Express i in a + ib form.
18 1
i
25. How many 4-letter codes can be formed using the first 10 letters of the English
alphabet, if no letter can be repeated?
m
c. o
26. Find an approximation of (0.99)5 using the first three terms of its expansion
m
co m
13
. e
27. The sum of first three terms of a G.P. is and their product is − 1. Find the
m las
12
s e
common ratio and the terms
g
l a a
ag
28. Find the equation of the ellipse whose vertices are (± 13, 0) and foci are (± 5, 0).
29. One card is drawn from a well shuffled deck of 52 cards. If each outcome is
equally likely, calculate the probability that the card will be (i) a diamond (ii)
not an ace card
PART –C
IV. Answer Any Six of the following questions (6 3 = 18)
m
30. Let U= {1, 2, 3, 4, 5, 6} A= {2, 3}, and B={3, 4, 5}. Show that (A ∪ B)∁ = A∁ ∩ B ∁
. co
3 1 . Determine the domain and range of the relation R defined by
em
R = {(x, x + 5): x ∈ {0, 1,2, 3, 4, 5}}.
l a s
32. Show that tan 3 x tan 2 x tan x = tan 3x – tan 2 x – tan x
a g
𝑢 𝑣
33. If (x + iy) 3 = u + iv, then show that 𝑥 + 𝑦 = 4(𝑥 2 − 𝑦 2 ).
34. Solve the inequality and show the graph of the solution on the number line
3x 4 x 1
-1
2 4
𝑥 2 4
35. Expand using binomial theorem (1 + 2 − 𝑥) , 𝑥 ≠ 0
36. If the angle between two lines is π/4 and slope of one of the lines is 1/2
find the slope of the other line.
m
c. o
37. If the origin is the centroid of the triangle PQR with vertices P (2a, 4, 6
o m m
c
Q(4,3b, 10) and R (8,14, 2c) then find the values of a, b, c
. s e
em l a
Find the derivative of 𝑡𝑎𝑛𝑥 w. r. t x from first principle.
s
38.
la a g
ag PART – D
V. Answer Any Four of the following questions (4 5 = 20)
39. Define Modulus function. Draw the graph of it. Also write its domain and
range.
cos 4 x cos 3 x cos 2 x
40. Prove that = cot 3x
sin 4 x sin 3 x sin 2 x
41. A committee of 7 has to be formed from 9 boys and 4 girls. In how many ways
m .
co
can this be done when the committee consists of:
(i) exactly 3 girls?
.
(ii) atleast 3 girls? (iii) atmost 3 girls?
e m
e m l a s
l a s a g
ag For more Question Papers, Sample Papers, Notes & Syllabus visit Page 3 of 4
Page 5
42. Derive a formula for the angle between two lines with slopes m1 and m2 . If the
1
angle between two lines is and slope of one of the lines is , find the slope of
4 2
the other line
𝑥 𝑛−𝑎 𝑛
43. Prove that,for any positive integer n,lim [ 𝑥−𝑎 ] = 𝑛𝑎𝑛−1 , and hence
𝑥→0
𝑥 15−1
evaluate lim [𝑥 10−1].
𝑥→1
44. Find the variance and standard deviation for the following data
xi 4 8 11 17 30 24 32
fi 3 5 9 5 4 3 1
45. One card is drawn from a well shuffled deck of 52 cards. If each outcome is
equally
likely, calculate the probability that the card will be
(i) a diamond (ii) not an ace card (iii) a black card
(iv) not a diamond (v) not a black card.
PART-E
VI. Answer the following question.
46. Prove geometrically that 𝑐𝑜𝑠(𝑥 + 𝑦) = 𝑐𝑜𝑠𝑥 𝑐𝑜𝑠𝑦 − 𝑠𝑖𝑛𝑥 𝑠𝑖𝑛𝑦, and hence
evaluate cos x
2
OR
𝑥2 𝑦2
Define hyperbola and derive the equation in the form 𝑎2 − 𝑏2 = 1 . 6
x5 cos x
47. Find the derivative of f x
sin x
OR
Find the sum of n terms of the sequence 7,77,777,7777,….tonterms. 4
*****************************************************************
For more Question Papers, Sample Papers, Notes & Syllabus visit Page 4 of 4